# Embedding in Recommender Systems: A Survey

### v2

@@ 3.3. Embedding of Heterogeneous Graphs
The heterogeneous graph is a more realistic representation of many real-world graphs than simplified graphs like homogeneous and bipartite graphs. These graphs encompass diverse node categories that can be interconnected. In contrast to homogeneous graphs with just one node type, heterogeneous graphs are more intricate and contain more abundant information. In the context of recommender systems, incorporating diverse relationships can lead to more accurate learned entity embeddings. For instance, combining a user-item bipartite graph with a social network of users can enhance the precision of predicting user preferences.
@@
Recent research focuses on enhancing user embeddings in recommender systems by incorporating social influence modeling. DiffNet (Wu et al., 2019a) and GraphRec (Fan et al., 2019b) tackle social and user-item networks separately. DiffNet oversimplifies social influence modeling by assuming uniform influence among a user’s neighbors. In contrast, GraphRec employs GAT (Veličković et al., 2017b) to better model real-world social influence by weighing friends’ influence based on the similarity of their initial embeddings. DiffNet++ (Wu et al., 2020a) builds upon DiffNet by introducing a unified framework that considers both social networks and user-item bipartite graphs. It addresses the limitation of DiffNet’s inability to distinguish varying degrees of influence from different neighbors through a multi-level attention mechanism. GES (Wang et al., 2018a) leverages related side information like brand to aggregate nodes’ embeddings for obtaining item embeddings in heterogeneous graphs. However, these methods primarily account for static social influence. DANSER (Wu et al., 2019b) introduces a dual GAT to capture both static and dynamic influence, acknowledging that a user’s impact on friends can vary based on items, leading to more realistic embeddings.
@@
The knowledge graph is also a type of heterogeneous graph that is widely used in recommender systems. Nodes in the knowledge graph represent entities, edges represent the relationships between entities, and edge attributes describe the nature of these relationships. By accurately describing real-world relationships, the knowledge graph provides rich semantic information that can improve interpretability and alleviate the sparsity problem in recommender systems. To learn entity embeddings in the knowledge graph, several translation-based models such as TransE (Bordes et al., 2013), TransH (Wang et al., 2014), and TransR (Lin et al., 2015) have been utilized. However, these approaches fail to fully utilize the topological and semantic information in the knowledge graph, which is precisely what GNNs excel at. To overcome this, methods like KGCN (Wang et al., 2019c) and KGAT (Wang et al., 2019a) apply GCN and GAT respectively to produce higher-quality embeddings. Nevertheless, these methods disregard users’ diverse intents when modeling user-item relationships, thus constraining embedding learning quality. Addressing this, KGIN (Wang et al., 2021a) introduces an intent layer between users and their interacted items for finer-grained relationship modeling.
@@ 3.4. Embedding of Hypergraphs
The hypergraph is a graph structure in which an edge can connect arbitrarily many nodes, and edges in a hypergraph are called hyperedges. In recommender systems, hypergraphs can be used to model complex entity relationships. A hypergraph can be represented as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathrm{H})$, where $\mathcal{V}$ represents the set of nodes, $\mathcal{E}$ represents the set of hyperedges, and $\mathrm{H}$ represents the incidence matrix.
@@
IHGNN (Cheng et al., 2022) enhances node embeddings through hypergraphs, revealing higher-order interaction patterns in user query histories. It introduces hyperedges $(u,q,p)$ for user query and product, employing three high-order feature interactions for hyperedge embedding aggregation. HyperGroup (Guo et al., 2021c) targets group recommendation, addressing potential misalignment between user and group preferences by representing groups as hyperedges. It aggregates neighboring hyperedges to capture group similarity. HEMR (La Gatta et al., 2022) focuses on music recommendation using hypergraph embeddings. It employs hyperedge-level random walks, followed by skip-gram for node embedding learning. The jump probability $p_{e}$ to each node, is defined as follows:
(13) $p_{e}=\min\left(\frac{\alpha}{D(e)}+\beta,1\right)$
Where $D(e)$ denotes the degree of the current hyperedge $e$, $\alpha$ and $\beta$ are hyperparameters. This approach encourages jumping out of low-degree hyperedges and exploring the complex relationships of nodes in the hypergraph.
The graph embedding methods mentioned above are summarized in Table 2. We classify the graph embedding methods according to the way they aggregate information in the graph. For example, random walk denotes aggregation of information through a sequence of nodes sampled by random walk.
3.5. Surveys & Tools
Surveys. Unlike our discussion focusing on learning graph embedding in recommender systems, Wang et al. (Wang et al., 2020a) focuses on the design of graph neural networks in different types of graph structures, while they do not consider the application of hypergraphs in recommender systems. Gao et al. (Gao et al., 2021) focuses on the application of graph neural networks in different recommender scenarios such as sequence recommendation and multi-behavior recommendation.
Tools. We also collected some high quality tools (Wang et al., 2020a) [^7] https://github.com/maenzhier/GRecX and repositories [^8] https://github.com/tsinghua-fib-lab/GNN-Recommender-Systems [^9] https://github.com/DLUTElvis/GNN4Rec-Papers on graph-based recommendation systems.
Table 2. Summary of Graph Embedding Methods
Graph type Method Message Passing
Homogeneous Graph DeepWalk (Perozzi et al., 2014), APP (Zhou et al., 2017), InfoWalk (Zhou et al., 2021) Random walk
M2GRL (Wang et al., 2020b), SGNS (Levy and Goldberg, 2014), DG-ENN (Guo et al., 2021b) Multiple view graphs + GCN
LINE (Tang et al., 2015) GCN
PGE (Li et al., 2017) Time decay weighted
Bipartite Graph GC-MC (Berg et al., 2017), STAR-GCN (Zhang et al., 2019b), NGCF (Wang et al., 2019b) Average neighbor embedding
LightGCN (He et al., 2020) Normalized summation
CSE (Chen et al., 2019), FBNE (Chen et al., 2020c) Random Walk
GE (Xie et al., 2016) Multiple view graphs + LINE
Heterogeneous Graph DiffNet (Wu et al., 2019a) Average neighbor embedding
GraphRec (Fan et al., 2019b), DiffNet++ (Wu et al., 2020a), DANSER (Wu et al., 2019b) GAT
GES (Wang et al., 2018a) GAT+Side Information
TransGRec (Wu et al., 2020b) GCN
GHL (Niu et al., 2020) Gated GNN
Hyper Graph IHGNN (Cheng et al., 2022), HyperGroup (Guo et al., 2021c) Weighted average of Hyperedge
HEMR (La Gatta et al., 2022) Hyperedge random walk
3.6. Future Direction
@@
Dynamic Graph Embedding for Recommendation. In graph-based recommender systems, dynamic graph evolution is vital. Embeddings from static graphs can result in uninteresting recommendations for users. Current research has rarely tackled dynamic graph embedding in recommender systems. While PGE (Li et al., 2017) considers temporal decay for item weights, it still treats graphs as static and overlooks evolving network structures. Leveraging dynamic graph neural networks (Manessi et al., 2020; You et al., 2022) for effective recommender system applications is a critical area for future exploration.
Fair Graph Embedding for Recommendation. Fair machine learning has recently emerged as a prominent area of focus in the machine learning community. Its primary objective is to alleviate bias of various forms in models developed through data-driven techniques. Within the domain of graph recommendation systems, the emphasis is on preventing models from relying on biased signals as patterns and subsequently propagating these biases throughout the graph. Given the profound societal impact of recommendation systems, it is crucial to ensure the acquisition of fair embeddings. While prior studies (e.g., GEAR (Ma et al., 2022) and FairGo (Wu et al., 2021a)) can apply to this, fair graph embedding remains underexplored. Also, benchmarking fairness in graph-based recommendation systems is a key research avenue. For a comprehensive understanding of fair machine learning, you can refer to the survey (Mehrabi et al., 2021).
Edge Embedding for Recommendation. Existing generic GNN frameworks seldom place a strong emphasis on handling edge features. GNNs that do incorporate edge features often adopt domain-specific strategies, customizing their approaches to address particular tasks. Examples of such tasks include quantum chemistry (Gilmer et al., 2017), circuit design (Zhang et al., 2019a), and graph matching (Yu et al., 2019). However, this specialized focus has contributed to a notable gap in research when it comes to exploring the significance of edge features within GNN-based recommender systems. Within the context of recommender systems, the majority of research efforts have traditionally centered on node features, user-item interactions, and graph structures, largely overlooking the potential wealth of information that edge features can provide. This oversight is surprising given that edges in recommender systems can encode essential information, such as the strength of user-item connections, the temporal dynamics of interactions, or the trustworthiness of user reviews. Neglecting edge features in the context of recommender systems not only limits our understanding of the nuances within user-item interactions but also hampers the development of more accurate and context-aware recommendation models. To address this gap, there is a growing need for research that delves into the role of edge features within GNN-based recommender systems. Investigating how to effectively incorporate and leverage edge features, alongside node features and graph structures.
@@ 4. Hash Embedding
In recommender systems, an alternative to the prevalent one-hot encoding for representing categorical attributes is hashing. Hashing embeddings were introduced to address the limitations associated with one-hot encoding. One-hot encoding often results in high-dimensional and extremely sparse feature matrices, posing significant scalability challenges, especially within complex deep recommendation systems. Hashing embeddings provide an effective solution to these challenges, offering benefits for both scalability enhancement and complexity reduction. Hashing embedding involves the application of one or more hash functions to the original sparse encoding, effectively curbing storage demands and computational overhead. Compared to raw one-hot encodings, hash embeddings manage to maintain recommendation effectiveness while substantially reducing the complexity of training and inference.
Considering various hash embedding methods based on factors like the number of hash functions, their utilization, and post-processing steps, we categorize the literature into three groups: 1) single function hash embedding; 2) multiple functions hash embedding; 3) dense hash embedding.
@@ 4.1. Single Function Hash Embedding
The hashing trick, initially introduced by Yahoo (Weinberger et al., 2009), uses a single hash function to map the features into hash values. This method is particularly relevant in collaborative filtering scenarios, exemplified by matrix factorization ($\mathbf{M}=\mathbf{U}\mathbf{V}^{T}\in\mathbb{R}^{m\times n}$), where $\mathbf{U}\in\mathbb{R}^{m\times d}$ and $\mathbf{V}$ $\in\mathbb{R}^{n\times d}$ are factorized matrices of the sparse $\mathbf{M}$. The hashing trick diminishes the dimensions of these matrices through a binary encoding function $f_{E}$ and hash function $f_{H}$. In this context, with feature index $(j,k)$, and position index $p$, $f_{E}$ transforms $(j,k)$ into values in $\{0,1\}$, effectively positioning the feature in hash table position $i$ using hash function $f_{H}$. This leads to $\mathbf{U^{\prime}}$in $\mathbb{R}^{{m^{\prime}}\times d}$ and $\mathbf{V^{\prime}}$in $\mathbb{R}^{n^{\prime}\times d}$, where $m^{\prime}\ll m$ and $n^{\prime}\ll n$, as follows:
[…]
@@
(16) $h_{ij}=\left\{\begin{aligned} &1,\ \rm{if}\;j\ \rm{mod}\ m=i\\ &0,\ \rm{otherwise}\\ \end{aligned}\right.$
The hashing trick has profoundly impacted the field of hash embedding within recommendation systems. This approach stands as the pioneer of simple, yet powerful techniques that markedly enhance the efficiency of original one-hot embeddings, particularly when dealing with high-dimensional features.
@@ 4.2. Multiple Functions Hash Embedding
Hybrid hashing (Zhang et al., 2020), introduced by Twitter, follows a similar premise. It employs two hash functions but applies them exclusively to infrequent feature values. For frequent feature values, the one-hot full embedding method is employed. This approach divides the embedding space considering feature frequencies, ensuring that vital information from the top-$K$ frequent features is retained. By employing a combination of one-hot full embedding and hybrid hashing, Facebook’s Q-R trick (Shi et al., 2020) adopts the complementary partition idea with multiple hash functions to avert hash collisions. Q-R trick introduces complementary hash functions through two embedding tables: $\mathbf{W}_{1}\in\mathbb{R}^{l\times d}$ and $\mathbf{W}_{2}\in\mathbb{R}^{\frac{a}{l}\times d}$ for each categorical feature with $a$ distinct values. The final embedding results from pooling the outputs of each hash function. The hash embedding combines two components: $\mathbf{W}_{1}$ represents the coarse-grained aspect where multiple values can map to the same row, while $\mathbf{W}_{2}$ functions as the fine-grained counterpart, compensating for collisions in $\mathbf{W}_{1}$. The ultimate hash embedding is computed as:
@@
Here, $\mathbf{H}$ and $\mathbf{H}^{\prime}$ denote corresponding hash matrices, and $\odot$ signifies element-wise multiplication. This design employing complementary hash functions ensures representation uniqueness.

### v3

@@ 4.4. Embedding of Heterogeneous Graphs
The heterogeneous graph is a more realistic representation of many real-world graphs than simplified graphs like homogeneous and bipartite graphs. These graphs encompass diverse node categories that can be interconnected. In contrast to homogeneous graphs with just one node type, heterogeneous graphs are more intricate and contain more abundant information. In the context of recommender systems, incorporating diverse relationships can lead to more accurate learned entity embeddings. For instance, combining a user-item bipartite graph with a social network of users can enhance the precision of predicting user preferences. Recent research focuses on enhancing user embeddings in recommender systems by incorporating social influence modeling. DiffNet (Wu et al., 2019a) and GraphRec (Fan et al., 2019) tackle social and user-item networks separately. DiffNet oversimplifies social influence modeling by assuming uniform influence among a user’s neighbors. In contrast, GraphRec employs GAT (Veličković et al., 2017b) to better model real-world social influence by weighing friends’ influence based on the similarity of their initial embeddings. DiffNet++ (Wu et al., 2020) builds upon DiffNet by introducing a unified framework that considers both social networks and user-item bipartite graphs. It addresses the limitation of DiffNet’s inability to distinguish varying degrees of influence from different neighbors through a multi-level attention mechanism. GES (Wang et al., 2018a) leverages related side information, like brand, to aggregate nodes’ embeddings for obtaining item embeddings in heterogeneous graphs. However, these methods primarily account for static social influence. DANSER (Wu et al., 2019c) introduces a dual GAT to capture both static and dynamic influence, acknowledging that a user’s impact on friends can vary based on items, leading to more realistic embeddings.
@@
The knowledge graph is also a type of heterogeneous graph that is widely used in recommender systems. Nodes in the knowledge graph represent entities, edges represent the relationships between entities, and edge attributes describe the nature of these relationships. By accurately describing real-world relationships, the knowledge graph provides rich semantic information that can improve interpretability and alleviate the sparsity problem in recommender systems. To learn entity embeddings in the knowledge graph, several translation-based models such as TransE (Bordes et al., 2013), TransH (Wang et al., 2014), and TransR (Lin et al., 2015) have been utilized. However, these approaches fail to fully utilize the topological and semantic information in the knowledge graph, which is precisely what GNNs excel at. To overcome this, methods like KGCN (Wang et al., 2019c) and KGAT (Wang et al., 2019a) apply GCN and GAT, respectively, to produce higher-quality embeddings. Nevertheless, these methods disregard users’ diverse intents when modeling user-item relationships, thus constraining embedding learning quality. Addressing this, KGIN (Wang et al., 2021a) introduces an intent layer between users and their interacted items for finer-grained relationship modeling.
@@ 4.5. Embedding of Hypergraphs
The hypergraph is a graph structure in which an edge can connect arbitrarily many nodes, and edges in a hypergraph are called hyperedges. In recommender systems, hypergraphs can be used to model complex entity relationships. IHGNN (Cheng et al., 2022) enhances node embeddings through hypergraphs, revealing higher-order interaction patterns in user query histories. It introduces hyperedges $(u,q,p)$ for user query and product, employing three high-order feature interactions for hyperedge embedding aggregation. HyperGroup (Guo et al., 2021c) targets group recommendation, addressing potential misalignment between user and group preferences by representing groups as hyperedges. It aggregates neighboring hyperedges to capture group similarity. HEMR (La Gatta et al., 2022) focuses on music recommendation using hypergraph embeddings. It employs hyperedge-level random walks, followed by skip-gram for node embedding learning.
4.6. Surveys and Future Directions
@@
Surveys. Unlike our discussion focusing on learning graph embedding in recommender systems, Wang et al. (Wang et al., 2020a) focus on the design of graph neural networks in different types of graph structures, while they do not consider the application of hypergraphs in recommender systems. Gao et al. (Gao et al., 2021) focus on the application of graph neural networks in different recommender scenarios such as sequence recommendation and multi-behavior recommendation. Embeddings from static graphs can result in uninteresting recommendations for users. While PGE (Li et al., 2017a) considers temporal decay for item weights, it still treats graphs as static and overlooks evolving network structures. Leveraging dynamic graph neural networks (Manessi et al., 2020; You et al., 2022) for effective recommender system applications is a critical area for future exploration.
Future Direction 1: Fair Graph Embedding for Recommendation. The pursuit of fairness in machine learning is gaining considerable traction, with a critical goal of mitigating biases inherent in data-driven models. In the context of graph-based recommender systems, preventing the learning and propagation of biased patterns is especially crucial due to their significant societal impact. Thus, ensuring the fairness of learned embeddings is a pressing concern. While initial studies in fair graph learning (e.g., GEAR (Ma et al., 2022) for fair graph classification and FairGo (Wu et al., 2021a) for fair influence maximization) provide relevant foundations, the specific area of fair graph embedding tailored for recommendation systems remains relatively underexplored. Developing robust methods to achieve fairness in these embeddings and establishing comprehensive benchmarks to evaluate fairness in graph-based recommendation systems are key research avenues. For a broader understanding of fair machine learning, readers are referred to the survey by Mehrabi et al. (Mehrabi et al., 2021).
Future Direction 2: Leveraging Topological Insights in GNNs for Enhanced Recommender Systems. Recent breakthroughs in topology-inspired graph neural networks (Papillon et al., 2023) present exciting opportunities for advancing recommender systems. For instance, curvature-based graph rewriting techniques have demonstrated success in mitigating common GNN pitfalls like over-smoothing and over-squashing (Topping et al., 2022; Nguyen et al., 2023; Fesser and Weber, 2024). Applying these techniques could potentially lead to more nuanced and informative representations of user-item interactions, especially in complex recommendation scenarios. Persistent Homology (PH) (Edelsbrunner et al., 2008), a powerful tool from topological data analysis for capturing multi-scale topological features, has been effectively integrated with GNNs. This integration has shown promising results and, in some cases, significant improvements in diverse graph-based tasks, including node classification (Horn et al., 2021), link prediction (Yan et al., 2021a), and the regression of molecular properties (Immonen et al., 2024). In recommender systems, applying PH to user-item interaction graphs could potentially uncover topological features (e.g., persistent cycles or connected components) that remain stable across various similarity thresholds. These features might correspond to robust patterns in user preferences or item relationships, offering novel insights to enhance collaborative filtering algorithms. Furthermore, the persistence diagrams generated through PH analysis could serve as unique topological signatures for users or items, thereby potentially enriching their feature representations in recommendation models. Additionally, advanced structural coding schemes (Zhao et al., 2021a; Zhang et al., 2023b), effective in identifying critical components within graphs, could be adapted to pinpoint influential users or pivotal items in recommendation networks. Such an adaptation could improve the precision of social and collaborative filtering recommendations.
@@ 5. Hash Embedding
In recommender systems, an alternative to the prevalent one-hot encoding for representing categorical attributes is hashing. Hashing embeddings were introduced to address the limitations associated with one-hot encoding. One-hot encoding often results in high-dimensional and extremely sparse feature matrices, posing significant scalability challenges, especially within complex deep recommender systems. Hashing embeddings provides an effective solution to these challenges, offering benefits for both scalability enhancement and complexity reduction. Hashing embedding involves the application of one or more hash functions to the original sparse encoding, effectively curbing storage demands and computational overhead. Compared to raw one-hot encodings, hash embeddings manage to maintain recommendation effectiveness while substantially reducing the complexity of training and inference.
Table 5. Hash Embedding
Hash Embedding Single Function Hash Embedding (§5.1) The Hashing Trick: (Yahoo) (Weinberger et al., 2009)
Multiple Functions Hash Embedding (§5.2) Key Models: Bloom Filter based (Bonomi et al., 2006), Hash Embedding (Tito Svenstrup et al., 2017), Hybrid Hashing (Twitter) (Zhang et al., 2020b), Q-R Trick (Facebook) (Shi et al., 2020)
DHE: Dense Hash Embeddings (§5.3) Key Models: Dense Hash Embedding (DHE) (Kang et al., 2020a), Dynamic Sparse Learning (DSL) (Wang et al., 2024b)
Considering various hash embedding methods based on factors like the number of hash functions, their utilization, and post-processing steps, we categorize the literature into three groups: 1) single function hash embedding; 2) multiple functions hash embedding; 3) dense hash embedding. A comparison of these categories is presented in Table 5.
@@ 5.1. Single Function Hash Embedding
The hashing trick, proposed by Yahoo (Weinberger et al., 2009), uses a single hash function to map features to hash values. This method is particularly relevant in collaborative filtering scenarios, exemplified by matrix factorization ($\mathbf{M}=\mathbf{U}\mathbf{V}^{T}\in\mathbb{R}^{m\times n}$), where $\mathbf{U}\in\mathbb{R}^{m\times d}$ and $\mathbf{V}$ $\in\mathbb{R}^{n\times d}$ are factorized matrices of the sparse $\mathbf{M}$. The hashing trick diminishes the dimensions of these matrices through a binary encoding function $f_{E}$ and a hash function $f_{H}$. In this context, with feature index $(j,k)$, and position index $p$, $f_{E}$ transforms $(j,k)$ into values in $\{0,1\}$, effectively positioning the feature in hash table position $i$ using hash function $f_{H}$. This leads to $\mathbf{U^{\prime}}$ $\in$ $\mathbb{R}^{{m^{\prime}}\times d}$ and $\mathbf{V^{\prime}}$ $\in$ $\mathbb{R}^{n^{\prime}\times d}$, where $m^{\prime}\ll m$ and $n^{\prime}\ll n$, as follows:
(14) $\left\{\begin{aligned} &\mathbf{U^{\prime}}_{pk}=\sum_{j:f_{H}(j,k)=p}f_{E}(j,k)\mathbf{U}_{jk}\\ &\mathbf{V^{\prime}}_{pk}=\sum_{j:f_{H}^{\prime}(j,k)=p}f_{E}^{\prime}(j,k)\mathbf{V}_{jk}\end{aligned}\right$.
@@
[…]
The hashing trick has had a profound impact on the field of hash embedding for recommendation systems. This approach stands as the pioneer of simple yet powerful techniques that markedly enhance the efficiency of original one-hot embeddings, particularly when dealing with high-dimensional features.
The primary challenge of this ”hashing trick” is the potential for hash collisions, where different features are mapped to the same embedding index, leading to information loss. To counteract this, multi-hash approaches have been proposed, as illustrated in Fig. 5.
Figure 5. The mechanism and challenge of hash embeddings. (Left) The ”hashing trick” maps high-cardinality IDs to a smaller embedding table. (Top-Right) A key challenge is hash collision, where different IDs (e.g., 897 and 2045) collapse to the same index, causing representation ambiguity. (Bottom-Right) Multi-hash methods provide a solution by using multiple hash functions to ensure unique representations.
@@ 5.2. Multiple Functions Hash Embedding
Hybrid hashing (Zhang et al., 2020b), introduced by Twitter, follows a similar premise. It employs two hash functions but applies them exclusively to infrequent feature values. For frequent feature values, the one-hot full embedding method is employed. This approach divides the embedding space considering feature frequencies, ensuring that vital information from the top-$K$ frequent features is retained. By employing a combination of one-hot full embedding and hybrid hashing, Facebook’s Q-R trick (Shi et al., 2020) adopts the complementary partition idea with multiple hash functions to avert hash collisions. Q-R trick introduces complementary hash functions through two embedding tables: $\mathbf{W}_{1}\in\mathbb{R}^{l\times d}$ and $\mathbf{W}_{2}\in\mathbb{R}^{\frac{a}{l}\times d}$ for each categorical feature with $a$ distinct values. The final embedding results from pooling the outputs of each hash function. The hash embedding combines two components: $\mathbf{W}_{1}$ represents the coarse-grained aspect, where multiple values can map to the same row, while $\mathbf{W}_{2}$ functions as the fine-grained counterpart, compensating for collisions in $\mathbf{W}_{1}$. The ultimate hash embedding is computed as:
@@
Here, $\mathbf{H}$ and $\mathbf{H}^{\prime}$ are the corresponding hash matrices, and $\odot$ is element-wise product. This design employs complementary hash functions to ensure representation uniqueness.
Recent comprehensive benchmarking studies have provided new insights into the effectiveness of hash embedding methods in real-world deployment scenarios. Tran et al. (Tran et al., 2025) conducted an extensive evaluation of lightweight embedding-based recommender systems, revealing that the performance of hash embedding methods varies significantly across different recommendation tasks. Their analysis shows that while hash embedding approaches like the Q-R trick achieve substantial memory reduction, the trade-offs between compression ratio and recommendation quality are highly task-dependent. The benchmark results indicate that simple magnitude-based pruning can sometimes outperform complex hash embedding schemes, suggesting that the field may benefit from focusing on fundamental compression principles rather than increasingly sophisticated hashing strategies.